Find the local and absolute extreme values of the function on the given interval. Also

specify the intervals where function is increasing or decreasing

𝑓(π‘₯) = π‘₯ + 2π‘π‘œπ‘ π‘₯


1
Expert's answer
2022-01-10T16:38:17-0500
"f(x)=x+2\cos x"


Domain: "(-\infin, \infin)"

Find the fist derivative with respect to "x"

"f'(x)=(x+2\cos x)'=1-2\sin x"

Find the critical number(s)


"f'(x)=0=>1-2\sin x=0"

"x=(-1)^n\dfrac{\pi}{6}+\pi n, n\in \Z"

Critical numbers: "(-1)^n\dfrac{\pi}{6}+\pi n, n\in \Z"

Find the second derivative with respect to "x"

"f''(x)=(1-2\sin x)'=-\cos x"




"f''(\dfrac{\pi}{6}+2\pi n)=-\dfrac{\sqrt{3}}{2}<0"

"f''(\dfrac{5\pi}{6}+2\pi n)=\dfrac{\sqrt{3}}{2}>0"

i) If "x\in [-\pi, \pi]"


"f(-\pi)=-\pi+2\cos (-\pi)=-\pi-2"

"f(\pi)=\pi+2\cos (\pi)=\pi-2"

"f(\dfrac{\pi}{6})=\dfrac{\pi}{6}+\sqrt{3}"

"f(\dfrac{5\pi}{6})=\dfrac{5\pi}{6}-\sqrt{3}"

The function "f(x)" has a local maximum on "[-\pi, \pi]" with value of "\dfrac{\pi}{6}+\sqrt{3}" at "x=\dfrac{\pi}{6}."

The function "f(x)" has a local minimum on "[-\pi, \pi]" with value of "\dfrac{5\pi}{6}-\sqrt{3}" at "x=\dfrac{5\pi}{6}."

The function "f(x)" has the absolute maximum on "[-\pi, \pi]" with value of "\dfrac{\pi}{6}+\sqrt{3}" at "x=\dfrac{\pi}{6}."

The function "f(x)" has the absolute minimum on "[-\pi, \pi]" with value of "-\pi-2" at "x=-\pi."

The function "f(x)" increases on "(-\pi, \dfrac{\pi}{6})\cup(\dfrac{5\pi}{6}, \pi)."

The function "f(x)" decreases on "(\dfrac{\pi}{6}, \dfrac{5\pi}{6})."


ii)

The function "f(x)" increases on "(\dfrac{5\pi}{6}+2\pi n, \dfrac{13\pi}{6}+2\pi n), n\in \Z."

The function "f(x)" decreases on "(\dfrac{\pi}{6}+2\pi n, \dfrac{5\pi}{6}+2\pi n), n\in \Z."